John is 3 times as old as Mary now and the sum of their present ages is 52.Let Mary's age be m.
(a)Express John's present age in terms of m.
(b)Form an equation in term of m and solve it.
(c)Hence ,find the present age of John.


please let me know the answer please.....

Answer by Theo(13342)

john is 3 times as old as mary.
the sum of their present ages is 52.
let mary's age be represented by m.
let john's age be represented by j.

the sum of their ages is equal to 52.

the equation for that is j + m = 52

john is 3 times as old as mary.

the equation for that is j = 3m

in the equation of j + m = 52, replace j with 3m to get:

3m + m = 52.

combine like terms to get:

4m = 52.

solve for m to get:

m = 52/4 = 13

since j = 3m, that makes j = 39.

john is currently 39 years old and mary is 13.

they told you to represent mary's age by m, but they didn't tell you what variable to use to represent john's age, so you were free to use any variable name you wanted to.

i chose j.